# IFT-UAM/CSIC-25-21

### v2

@@ E.1 Collisionless regime computation
where $\Delta m^{2}$ has to be understood as the change of mass of the particle $i$ upon crossing the wall. By conservation of momentum, the wall receives an equal and opposite kick, $\Delta p^{\rm part}_{z}=-\Delta p^{\rm wall}_{z}>0$, which accelerates it forward or backwards depending on the sign of the kick. We observe that a particle gaining mass induces a negative kick, and so resists the expansion of the wall, while a particle losing mass sucks the wall. In the model under consideration, both types of particles are present, so the competition between them will determine the sign of the collisionless pressure. To capture the pressure induced by the plasma, we need to further convolute the momentum with the incoming flux:
[…]
where the sum is to be performed over all particles which may lose or gain a mass across the bubble wall, and $g_{i}$ is the number of dofs for each particle. By convention, a negative pressure sucks the wall while a positive one resists the expansion. The integral over the phase space is frame-independent and we compute it in the plasma frame. On the other hand, the force exerted by the vacuum energy is given by
[…]
@@ E.2 Local thermal equilibrium approach
[…]

### v3

@@ E.1 Collisionless regime computation
where $\Delta m^{2}$ has to be understood as the change of mass of the particle $i$ upon crossing the wall. By conservation of momentum, the wall receives an equal and opposite kick, $\Delta p^{\rm part}_{z}=-\Delta p^{\rm wall}_{z}>0$, which accelerates it forward or backwards depending on the sign of the kick. We observe that a particle gaining mass induces a negative kick, and so resists the expansion of the wall, while a particle losing mass aspires the wall. In the model under consideration, both types of particles are present, so the competition between them will determine the sign of the collisionless pressure. To capture the pressure induced by the plasma, we need to further convolute the momentum with the incoming flux
$\mathcal{P}_{\rm plasma}$ $=\int dz\partial_{z}\phi\sum_{i}g_{i}\frac{dm^{2}_{i}(\phi)}{d\phi}\int\frac{d^{3}{\bf p}}{(2\pi)^{3}2E_{i}}\,f_{i}(p,z,T)\approx\sum_{i}g_{i}\int\frac{d^{3}{\bf p}}{(2\pi)^{3}}\frac{\Delta m^{2}_{i}}{2E_{i}}f^{\rm eq}_{\rm outside}(p,T)\,.$ (29)
where the sum is to be performed over all particles which may lose or gain a mass across the bubble wall, and $g_{i}$ is the number of dofs for each particle. By convention, a negative pressure aspires the wall while a positive one resists the expansion. The integral over the phase space is frame-independent and we compute it in the plasma frame. On the other hand, the force exerted by the vacuum energy is given by
${}F_{\rm vacuum}\equiv\int dz\partial_{z}\phi\frac{dV^{T=0}(\phi)}{d\phi},$ (30)
@@ E.2 Local thermal equilibrium approach
$\mathcal{P}_{\rm plasma}\equiv-\int dz\partial_{z}\phi\sum_{i}\frac{dm^{2}_{i}(\phi)}{d\phi}\int\frac{d^{3}{\bf p}}{(2\pi)^{3}2E_{i}}\,f_{i}(p,z,T)=\mathcal{P}_{\rm LTE}+\mathcal{P}_{\rm dissipative}\ .$ (32)